At Halloween party, Fancy is throwing darts at jack-o-lanterns and witches. She gets 5 points for every jack-o-lantern she hits and 8 points for every witch. She has hit 13 jack-olanterns and witches altogether and has 80 points. How many jack-o-lanterns and how many witches has she hit?
Answer by stanbon(75887)

At Halloween party, Fancy is throwing darts at jack-o-lanterns and witches. She gets 5 points for every jack-o-lantern she hits and 8 points for every witch. She has hit 13 jack-olanterns and witches altogether and has 80 points. How many jack-o-lanterns and how many witches has she hit?
---
Hit Equation: j + w = 13
Points Equat:5j +8w = 80
-----------------------------
Multiply thru 1st Eq. by 5:
5j + 5w = 5*13
5j + 8w = 80
---------------------
Subtract 1st from end and solve for "w":
3w = 80-5*13
3w = 15
w = 5 (# of witches hit)
---
Since j+w = 13, w = 8 (# of jacko's hit)
=============================================
Cheers,
Stan H.